OpenStudy (anonymous):

A contractor has a daily payroll of $277 when employing some men at $8 per day and the others at $ 9 per day . On increasing those at $ 8 by 50% and doubling those at $ 9, he creates a daily payroll of $ 474 . Find how many he originally employed at each wage .

OpenStudy (anonymous):

please help! :(

ganeshie8 (ganeshie8):

say the contractor has \(a\) employees at $8 and \(b\) employees at $9 : \[8a+9b = 277\tag{1}\] \[8(1.5a)+9(2b)=474\tag{2}\] two equations and two unknonws, you can solve them

OpenStudy (anonymous):

how to solve the two unknowns?

OpenStudy (anonymous):

 look pls

OpenStudy (anonymous):

a = 474 - 18b/8.5

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

a=474-18*b/8.5

OpenStudy (anonymous):

a=474-2.11765 a=?

OpenStudy (anonymous):

where did you get the 2.11765?

OpenStudy (anonymous):

by calculator.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

did u get the answer.

OpenStudy (anonymous):

471.88235

OpenStudy (anonymous):

linear equation with 2 variables use elimination 8a+9b=277 euqation 1 8(1.5a)+9(2b)=474=12a+18b=474 equation 2 then (8a+9b=277)3 (12a+18b=474)-2 24a+27b=831 -24a-36=-948 then cancel out the a since its positive and negative -9b=-118 -9b/-9=-118/-9 b=13 you can now use the value of b to substitute it into the equation 2